OpenStudy (anonymous):

What is the solution of sqrt(5x+1)-sqrt(x)=5 A. x=0 B. x=16 and x=0 C. X=16 D. x=16 and x=1

OpenStudy (anonymous):

square both sides to eliminate the sqrt, then try to solve for x

OpenStudy (shamil98):

or just by looking at the answer choices, it's obviously not 0.. nor is it 16 and 0 or 16 and 1 so just 16..

OpenStudy (zzr0ck3r):

plug and play:)

OpenStudy (shamil98):

\[\huge \sqrt { 5(16) + 1} - \sqrt{16} = 5\] \[\huge 9 -4 = 5 ~~\checkmark\]

OpenStudy (anonymous):

Yep its 16 thanks everybody